Line BD bisects angle ABC, angle ABC = 9x. Angle ABD = 4x + 13. Find angle DBC
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39617)
Answer by ikleyn(52784)

.
From the condition, you have


the measure of the angle ABC = two times the measure of the angle ABD.


In terms of "x" it means


    9x = 2*(4x + 13)

    9x = 8x + 26

    x = 26  degrees.


Then the measure of the angle DBC = 4x + 13 = 4*26 + 13 = 117 degrees.


Answer.  The measure of the angle DBC is 117 degrees.

Solved.